OpenStudy (anonymous):

Solve for x: 2|x − 3| + 1 = 7 x = 0, x = 6 x = −1, x = 6 x = 0, x = 9 No solutions

OpenStudy (hoblos):

2|x − 3| + 1 = 7 2|x − 3| = 6 |x − 3| = 3 x-3 = 3 or x-3 = -3 x= 6 x=0 first answer

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

glad to help (: